01:07:50 Submit Quin 9 of 12 (8 complete) This Quiz: 12 pts A large philanthropic organization keep s records on the people who have contributed to their cause. The organization buys demographic data on neighborhoods. A histogram and right. Complete parts (a) through (e). statistics for the percentage of whites in the neighborhoods of 500 donors are shown to the Mean 83.14 Median S3 StdDev22.59 0 25 50 75 100 b) Which is a better summary of the spread, the IQR or the standard deviation? r summary of the spread because the a normal model, about what percenta ge of neighborhoods should have a percent B. 67% @) E. 95% C. 68% OF 99.7% OA, 68% OD, 81.4% d) What percentage of neighborhoods actually have a percent white within one O A. Between 75% and 80% standard O B. D. Between 20% and 25% Between 85% and 90% 。C. Between 65% and 70% e) Explain the problem in using the normal model for these data. Select all that apply B. The data are symmetric The distrbution is skewed to the nght C. A. The distribution is skewed to the left F. The data are not unimocal E. The data are not bimodal Click to select your answer(s).
Answer #1

b) The correct option is IQR in the first box of this part.

c) The correct option is C) 68%.

d) The correct option is D

Because here mean = 83.14 and standard deviation = 22.59

So that the 1\sigma lower and upper limits are as follows.

Lower limit = \mu -  1\sigma = 83.14 - 22.59 = 60.55

and upper limit =  \mu -  1\sigma = 83.14 + 22.59 = 105.73

From the given histogram most of the values in this range.